OpenStudy (anonymous):

One dimension of a rectangular map is 8 inches and the perimeter is 40 inches. What is the other dimension of the map? A. 14 inches B. 12 inches C. 4 inches D. 2 inches

OpenStudy (***[isuru]***):

 2x + 16 = 40 x = 24/2 = 12 got it :) ?

OpenStudy (anonymous):

waooo thanks broo....